# Learning from Negative Samples in Generative Biomedical Entity Linking

### v1

@@ Appendix C Top-5 Accuracy
Table C presents our model’s top-1 and top-5 accuracy on the BC5CDR and AAP datasets. It compares the performance of our model in its baseline form (GenBioEL) and after fine-tuning (ANGEL) and combined pre-training and fine-tuning (ANGEL). Our approach consistently boosts top-1 accuracy across all datasets, though the trends in top-5 accuracy are less uniform. In BC5CDR, both top-1 and top-5 accuracy show significant improvements: top-1 accuracy rises by 1.4 percentage points (from 93.1% to 94.5%), and top-5 accuracy increases by 1.1 percentage points (from 95.7% to 96.8%). However, the AAP dataset exhibits a different pattern. While top-1 accuracy improves by 0.9 percentage points (from 89.3% to 90.2%), top-5 accuracy slightly declines: there is a 0.7 percentage points drop (from 95.4% to 94.7%) after fine-tuning and a 0.2 percentage points decrease (from 95.4% to 95.2%) after combined pre-training and fine-tuning. This decline in top-5 accuracy may be due to the AAP dataset’s limited contextual information, forcing the model to rely predominantly on the mention form, making it more challenging to maintain high accuracy across multiple predictions. Additionally, the negative sampling strategy could unintentionally bias the model toward optimizing top-1 accuracy, thereby impacting top-5 performance.
In conclusion, while our method consistently improves top-1 accuracy, the occasional slight decreases in top-5 accuracy, as observed in the AAP dataset, underscore the need for further refinement to maintain balanced accuracy across different ranking levels. Future work should focus on training strategies that preserve or enhance top-5 accuracy alongside top-1 improvements.
Appendix D Hyperparameter Configurations
| Pre-training | Fine-tuning | |||||
| NCBI | BC5CDR | COMETA | AAP | MM-ST21pv | ||
| Positive-only Training | ||||||
| Training Steps | 80K | 20K | 30K | 40K | 30K | 40K |
| Learning Rate | 4e-5 | 3e-7 | 5e-6 | 5e-6 | 5e-6 | 2e-5 |
| Weight Decay | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 |
| Batch Size | 384 | 16 | 16 | 16 | 16 | 16 |
| Adam | 1e-8 | 1e-8 | 1e-8 | 1e-8 | 1e-8 | 1e-8 |
| Warmup Steps | 1,600 | 0 | 500 | 500 | 0 | 1,000 |
| Attention Dropout | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
| Clipping Grad | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
| Label Smoothing | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
| Negative-aware Training | ||||||
| Epochs | 5 | 1 | 1 | 1 | 1 | 1 |
| Learning Rate | 1e-5 | 1e-5 | 1e-5 | 5e-6 | 5e-6 | 5e-6 |
| (DPO) | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
| Weight Decay | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 |
| Batch Size | 1,024 | 64 | 16 | 64 | 8 | 64 |
| Warmup Steps | 100 | 100 | 100 | 100 | 100 | 100 |
Table D details the hyperparameters used for positive-only training and negative-aware training across the BioEL benchmark datasets. We search for the hyperparameter settings that are optimized for each dataset. We refer to the study of Yuan et al. (2022b) to determine the range of the hyperparameters. During pre-training, we use the same hyperparameters as in GenBioEL. For positive-only training, we explore a range of training steps between 20K and 40K, a learning rate between 2e-5 and 3e-7, and batch sizes from 8 to 16, except during pre-training. During negative-aware training, we fix the at 0.1, in accordance with the basic configuration of DPO, and search the hyperparameter space using a learning rate between 1e-5 and 1e-6 and batch sizes ranging from 8 to 64.

### v2

@@ Appendix E Top-5 Accuracy
Table E presents our model’s top-1 and top-5 accuracy on the BC5CDR and AAP datasets. It compares the performance of our model in its baseline form (GenBioEL) and after fine-tuning (ANGEL) and combined pre-training and fine-tuning (ANGEL). Our approach consistently boosts top-1 accuracy across all datasets, though the trends in top-5 accuracy are less uniform. In BC5CDR, both top-1 and top-5 accuracy show significant improvements: top-1 accuracy rises by 1.4 percentage points (from 93.1% to 94.5%), and top-5 accuracy increases by 1.1 percentage points (from 95.7% to 96.8%). However, the AAP dataset exhibits a different pattern. While top-1 accuracy improves by 0.9 percentage points (from 89.3% to 90.2%), top-5 accuracy slightly declines: there is a 0.7 percentage points drop (from 95.4% to 94.7%) after fine-tuning and a 0.2 percentage points decrease (from 95.4% to 95.2%) after combined pre-training and fine-tuning. This decline in top-5 accuracy may be due to the AAP dataset’s limited contextual information, forcing the model to rely predominantly on the mention form, making it more challenging to maintain high accuracy across multiple predictions. Additionally, the negative sampling strategy could unintentionally bias the model toward optimizing top-1 accuracy, thereby impacting top-5 performance. In conclusion, while our method consistently improves top-1 accuracy, the occasional slight decreases in top-5 accuracy, as observed in the AAP dataset, underscore the need for further refinement to maintain balanced accuracy across different ranking levels. Future work should focus on training strategies that preserve or enhance top-5 accuracy alongside top-1 improvements.